-5x-3y+6z=-26
X+2y-2z=8
5x-4y+3z=12

Answer by sachi(548)

-5x-3y+6z=-26.........1
X+2y-2z=8 ........2
5x-4y+3z=12 ...........3

adding eqn 1 & 2 X 5

7y-4z=14 .....4
Adding eqns 1 & 3
-7y+9z=-14 .........6
adding eqns 4& 5
5z=0
or z=0
from eqn 4
y=2
from eqn 1
x=4